How do you find the slope intercept equation of a line through the given point (-5,1) and the slope of -1
Answer by stanbon(75887)

How do you find the slope intercept equation of a line through the given point (-5,1) and the slope of -1
-----------------------
Form: y = mx+b
You know y = 1 when x = -5, and you know m = -1;
So solve for "b":
--------
1 = -1*-5 + b
1 = 5+b
b = -4
-----
Equation:
y = -x-4
========================
Cheers,
Stan H.
========================